OpenStudy (anonymous):

x-1 x+3 ------=----- x+1 x-4

OpenStudy (dumbcow):

cross - multiply (x-1)(x-4) = (x+1)(x+3) x^2 -5x +4 = x^2 +4x +3 -5x +4 = 4x+3 -9x = -1 x = 1/9